Identify all coterminal angles between -360 degrees and 360 degrees for the angle -420 degrees.
A. -60 degrees and 300 degrees
B. -30 degrees and 330 degrees
C. 30 degrees and -330 degrees
D. 60 degrees and -300 degrees
Thank you so much for all your help!!!!
God Bless and Happy Holidays!!!

Answer by stanbon(75887)

Identify all coterminal angles between -360 degrees and 360 degrees for the angle -420 degrees.
---
You can always add or subtract whole number multiples of 360.
---
-420+360 = -60
-420+2*360 = 300
---------------------------
Ans: A
==================
Cheers,
Stan H.
==================
A. -60 degrees and 300 degrees
B. -30 degrees and 330 degrees
C. 30 degrees and -330 degrees
D. 60 degrees and -300 degrees